# MUMU: Bootstrapping Multimodal Image Generation from Text-to-Image Data

### v1

@@ 1 Introduction
We construct a multimodal training set bootstrapped from text-image data. We use open vocab object detection to extract image crops corresponding to words in the image captions [22]. The image crops are then placed before their corresponding words in the text prompt, see Figure 3 for an example. Our data is mostly synthetically generated from SDXL with some added high quality publicly available data.
@@ 2.3 Natively Multimodal Models
Open VLMs usually hybridize two pre-trained backbones. Existing work has demonstrated the image generation capabilities of natively trained multimodal models [35, 36]. However, these models are closed weights and cannot be directly built upon. Our work shows that a multimodal image generation model can be constructed from existing open weights models allowing for community research on this topic.
@@ 7 Some Open Questions Beyond Scaling
token based decoder   We use a diffusion decoder because SDXL is the highest quality open source image model. However, MUMU is decoder agnostic and could be retrained with an autoregressive or other token-based image generation backbone [60, 61, 62, 63, 2, 64, 35, 36]. Unfortunately, most of these backbones are closed source, and open source token based models [65, 28] generate lower quality images than SDXL.

### v2

@@ 1 Introduction
We construct a multimodal training set bootstrapped from text-image data. We use open vocabulary object detection to extract image crops corresponding to words in the image captions [22]. The image crops are then placed before their corresponding words in the text prompt, see Figure 3 for an example. Our data is mostly synthetically generated from SDXL with some added high quality publicly available data.
@@ 2.3 Natively Multimodal Models
Open VLMs usually hybridize two pre-trained backbones. Existing work has demonstrated the image generation capabilities of natively trained multimodal models [35, 36, 37]. However, these models are closed weights and cannot be directly built upon. Our work shows that a multimodal image generation model can be constructed from existing open weights models allowing for community research on this topic.
@@ 2.4 Diffusion Model Image Conditioning
Prior work [46] has considered combining relatively small multimodal encoders with SD1.5 using a similar training idea to the one presented here (object detection and interleaving captions) along with an alignment and instruction training phase. We add to this work by showing that just the interleaved object pre-training phase along with scaling the encoder and decoder is enough to achieve high quality multimodal control. In addition, we argue that training the entire model end-to-end rather than freezing decoders yields better outcomes.
@@ 7 Some Open Questions Beyond Scaling
token based decoder   We use a diffusion decoder because SDXL is the highest quality open source image model. However, MUMU is decoder agnostic and could be retrained with an autoregressive or other token-based image generation backbone [62, 63, 64, 65, 2, 66, 35, 67, 37]. Unfortunately, most of these backbones are closed source, and open source token based models [68, 28] generate lower quality images than SDXL.